the length of a rectangle is 6 yards longer than its width the perimeter is 32 what is its area
Answer by JulietG(1812)

Perimeter = 2 length + 2 width
P = 32
L = W+6
Take the known value of L and put it into the initial equation
P = 2L + 2W
32 = 2(W+6) + 2W
32 = 2W + 12 + 2W
32 = 4W + 12
Subtract 12 from each side
20 = 4W
Divide each side by 4
5 = W
If the width is 5, then the length is 5+6, or 11
5+5+11+11=32